OpenStudy (zmudz):

Find all \(2 \times 2\) matrices \(A\) that have the property that for any \(2 \times 2\) matrix \(B\), \(A B = BA \). Please help! Thank you!

OpenStudy (anonymous):

I believe this would be symmetric matrices.

OpenStudy (anonymous):

Just a hunch

OpenStudy (anonymous):

Symmetric matric matricies follow the pattern:\[ A = \begin{bmatrix} a&b\\b&c \end{bmatrix} \]

jimthompson5910 (jim_thompson5910):

I'm no sure if this encapsulates all cases, but here is a condition where A*B = B*A http://math.stackexchange.com/questions/170241/when-is-matrix-multiplication-commutative

jimthompson5910 (jim_thompson5910):

the user `Johannes Kloos` gives a good proof

jimthompson5910 (jim_thompson5910):

sorry typo, I meant to say "not" instead of "no"

OpenStudy (kainui):

One way is to "brute force" it and write out two matrices and multiply them together in both ways and set the corresponding entries equal and see what pops out.